# machine learning, optical communications, passive optical network, equalization, optical performance monitoring, modulation format identification, bit error rate, optical signal-to-noise ratio, nonlinearities

### v1

@@ 5.3 Transformer-based Network
The transformer framework emerged as the predominant approach for sequential data analysis after 2017. This shift in dominance is primarily due to the transformer’s remarkable predictive accuracy demonstrated in empirical studies, as well as its exceptional ability to capture intricate dependencies within long sequences VaswaniA (2017). While the transformer was initially designed for natural language processing tasks, it remains the most extensively utilized sequence model and serves as the fundamental framework for numerous models WuH (2022). However, in this section, we will comprehensively elucidate the design rationale behind each module. The overview of the time series transformer is illustrated in Fig. 2.
Figure 2: Transformer Architecture
Encoder-Decoder: The objective of an encoder-decoder is to discover a reduced-dimensional representation of the input at an intermediate layer and subsequently reconstruct the output with minimal error. By doing so, the encoder-decoder acquires the ability to compress and reconstruct the input data nonlinearly. The transformer model’s encoder component consists of N stacked sub-encoders. Each sub-encoder encompasses a self-attention layer, a feed-forward layer, a residual connection, and layer normalization. The initial sub-encoder processes the input embeddings vector, while subsequent sub-encoders take the output from the previous sub-encoder as their input.
Attention: The Scaled Dot-Product Attention mechanism is the key component aiming to aggregate information across different parts of the input sequence. Each input vector is transformed into three distinct vectors: Queries ($Q$), Keys ($K$), and Values ($V$). The process involves calculating the dot products of the queries with all keys, scaling them by the square root of the dimension dk, and applying a softmax function to obtain weights on the values VaswaniA (2017).
@@
Transformers employ multiple attention heads to implement multiple perspectives for different parts of the input sequence, allowing the model to simultaneously learn and consider multiple perspectives on the input data. Equation 14 and 15 represent the functionality of the multi-head attention step. $head_{i}$ represents the single attention head. The final result of the multi-head attention is concatenating all the attention heads.
@@
Although the vanilla transformer has demonstrated remarkable performance in dealing with time series data, the computational complexity of its original attention mechanism hampers its ability to process long sequences effectively. To overcome this limitation, researchers have developed various attention mechanism variants. An example of such is the Locality-Sensitive Hashing (LSH) Attention mechanism, which was introduced as part of the Reformer architecture KitaevN (2020). The LSH Attention mechanism utilizes specialized hash functions to transform queries and keys, thereby facilitating the categorization of similar items into shared hash buckets. Through sorting tokens based on their hash codes, items with similarities are grouped together, enabling the aggregation of relevant information. To enable parallel processing, the sorted sequence is divided into chunks. Subsequently, attention mechanisms are selectively applied to these chunks and their neighboring segments, allowing for focused examination of localized portions. By employing a systematic framework based on hash coding, the LSH Attention mechanism significantly improves computational efficiency compared to the vanilla transformer. It effectively addresses the challenge of processing long sequences by reducing computational complexity while preserving performance.
@@
Decoder: The decoder consists of several stacked sub-decoders. The ground truth follows a process similar to that of the encoder, being transformed into Query $Q$, Key $K$, and Value $V$ representations. Attention weights are calculated by comparing the Queries $Q$ from the decoder with every value $V$ in the encoder. This process is repeated in parallel across N sub-decoders, resulting in a final attention matrix. The attention matrix then undergoes a softmax operation, yielding probabilities for each value. In addition to the two sub-layers in each encoder layer, the decoder introduces a third sub-layer, performing multi-head attention over the encoder stack’s output VaswaniA (2017).
@@
Throughout the decoding process, the model is auto-regressive, taking previously generated outputs as additional input for developing the following output. Residual connections and layer normalization are applied around each sub-layer, and a specific self-attention sub-layer modification is made to avoid positions from attending to subsequent positions. Combined with the offset output embeddings, this ensures that predictions for a given position depend only on known outputs at positions earlier in the sequenceVaswaniA (2017).
@@
Attention Variants: The objective of the attention layer is to identify inter-dependencies and connections among various input embeddings, allowing the model to evaluate the importance of each element in relation to others. It explicitly computes relationships between different elements in the sequence, which can be insightful and intuitive for understanding how information flows through the model. Except for the problem of computational complexity, the mechanism in vanilla transformer VaswaniA (2017) still needs to be improved in processing inter-dependencies and periodicity of signal data. A new type of encoder in the Autoformer WuH (2021) replaces the original encoder, which applies series decomposition and autocorrelation to detect inter-dependencies, combining representations within the sub-series. The series decomposition component divides the original signal into two distinct parts: the seasonal component, which captures short-term patterns, and the trend component, which captures long-term behavior. This partitioning allows for identifying and representing both the short-term and long-term characteristics present in the time series data. Additionally, the auto-correlation mechanism utilizes Fast Fourier transforms (FFT) to compute correlations between the time series and its delayed version, providing insights into how the series relates to its past values at different time lags. The combination of series decomposition and auto-correlation effectively captures and represents the underlying trend and seasonality within the time series data.
@@
Recent Progress: Within this section, we provide a comprehensive overview of the vanilla transformer and its module design, particularly within the domain of time series and traffic prediction. Over the years, substantial improvements have been made to enhance the transformer for accurate time series prediction. Notably, advancements have been achieved in reducing computational complexity while improving the effectiveness of the attention mechanism logsparse (2019) informer (2021). However, recent research has introduced compact models based on multi-scale transformation scaleformer (2023), which unexpectedly surpass the performance of existing meticulously designed models across various benchmarks. This development has sparked a renewed and significant debate on the fundamental structure of sequence models. In the subsequent sections, we will summarize and delve deeper into this particular model, shedding light on its intricacies and implications. For the latest work, please refer to Table 3.
@@ 5.4 Fourier Convolution Neural Network
In the previous section, we introduced models based on convolutional kernels and subsampling as fundamental modules. Their core principle involves decomposing signals into different scales in the time domain and subsequently applying nonlinear transformations to learn salient features. However, for the majority of real-world signals, transforming them into the Fourier domain is often more efficient. This efficiency is attributed to the following factors: 1) The majority of real-world signals are bandpass or lowpass, and in the Fourier domain, their dynamic range decreases from $n$ to $exp(-n)$; 2) Fourier transformation is a bijective transformation, ensuring energy conservation and controllable error in both forward and inverse transformations; 3) The computational complexity of existing fast Fourier transform algorithms, after improvements, is $O(nlog(n))$, making it convenient for hardware implementation.
In this section, we will introduce TimesNet WuH (2022), while TimesNet utilizes a frequency-attention mechanism and FreTS YiK (2023) explicitly performs non-linear transformations in the frequency domain. For extensive models please refer to Table 3. The fundamental concept of TimesNet involves transforming the initial signal into $k$ distinct 2D tensors instead of directly processing the original sequence. This approach empowers the model to effectively capture both intraperiodic and interperiodic variations within these fixed windows. A variant based this model has achieved ShaoFConv (2024)
@@
Component Detection: Identifying the most crucial k frequencies is based on the amplitude of Fourier coefficients. Subsequently, utilizing the selected components within the k frequency range, the signal undergoes sampling using a continuous sampling method, and these subseries are aligned into a two dimensional tensor.
Alignment: The aligned subseries are then fed into a convolution-based module, specifically Inception, to alleviate distortion caused by channel effects. Fourier coefficients pass through a softmax function to generate attention weights, which are multiplied by the output of each convolution module to produce the final output.
Fourier Attention: Fourier transformation, being a global operation, induces periodic oscillations throughout the entire signal when there is any alteration in amplitude. Substantial variations in the amplitude of the primary components result in significant fluctuations. Leveraging this characteristic, TimesNet utilizes the Fourier spectrum of the nonlinearly transformed signal to determine the attention value for each component.
@@
FreTS: As the local characteristics vary, any time-domain-based processing is constrained by the limitations imposed by information bottlenecks. In comparison, FreTS explicitly utilizes frequency-domain features to directly mitigate distortion without directly manipulating the time-domain in its model architecture. FreTS is an MLP-based network that is able to effectively learn patterns of time series data in the frequency domain. As presented in YiK (2023), FreTS consists of two learners: Frequency Channel Learner and Frequency Temporal Learner. In fact, in the equalization problem, we do not have a so-called channel dimension but rather a stack of independent experiments. Thus, we only introduce a frequency domain MLP.
Frequency MLP: The frequency temporal learner aims to capture temporal patterns in the frequency domain. Specifically, for a complex number input $\mathcal{H}\in\mathbb{C}^{m\times d}$, the MLP aims to optimize the weight matrix $\mathcal{W}\in\mathbb{C}^{d\times d}$ and bias $\mathcal{B}\in\mathbb{C}^{d}$ so that final output $\mathcal{Y}\in\mathbb{C}^{m\times d}$ could approximately reconstruct the ground truth.
@@
The MLP in the frequency domain implies the global convolutions in the time domain. The detailed proof can be found in YiK (2023). An increasing work has demonstrated the feasibility of deep learning models in the frequency domain. Simultaneously, the corresponding computational complexity of frequency-domain processing has decreased from $O(n)$ to $O(nlogn)$ due to the reduction in the signal’s dynamic range. However, the advantages and disadvantages of networks in both the frequency and time domains remain inadequately explored. Due to space limitations, we offer a more detailed categorization, along with corresponding references and keywords, in Table 3 for researchers with specific interests.
@@ 6 Model Compression
In recent years, the proliferation of large-scale ML models has significantly advanced state-of-the-art technology across various domains, ranging from natural language processing to computer vision. The surge in model complexity, often characterized by intricate architectures and many parameters, has driven the need for efficient hardware implementations to harness their full potential. The advent of single graphics processing units (GPUs) as a critical computational resource has been pivotal in this landscape, offering a parallelized architecture suitable for accelerating the training and inference processes LewJ (2019). The significance of deploying large machine-learning models on a single GPU lies in optimizing computational efficiency and reducing latency. Single GPU implementations facilitate parallel processing, enabling the simultaneous execution of multiple tasks and handling extensive model parameters. This enhances the speed of model training and expedites real-time inference, a critical requirement in applications such as autonomous systems and edge computing. However, the hardware implementation of large machine-learning models on a single GPU is challenging. The complexity of these models often exceeds the computational capacity and memory constraints of a single GPU, necessitating innovative solutions for efficient utilization MarculescuD (2018). Techniques such as model pruning, quantization, vector quantization, and knowledge distillation have emerged as strategies to mitigate these challenges, ensuring that even formidable models can be accommodated within the limitations of a single GPU without compromising performance. The authors examine how to use a single GPU effectively for implementing large machine learning models. They discuss methods that balance complexity and computational efficiency to maximize hardware utilizationMarculescuD (2018).
@@
However, conducting a comprehensive performance-versus-complexity analysis is required to evaluate the suitability of various ANNs in short-reach optical communication systems. Deep learning models, including convolutional neural networks (CNNs), recurrent neural networks (RNNs), and long short-term memory networks (LSTMs), find applications in critical tasks such as equalization, fault detection, subcarrier allocation, nonlinearity compensation, and bandwidth request and allocation. The complexity of these models is a significant factor affecting their feasibility and use in practice. For instance, CNNs, renowned for their effectiveness in image processing, may introduce convolutional and pooling layers, increasing model complexity. Similarly, RNNs and LSTMs, designed for sequential data, introduce recurrent connections that enhance their ability to capture temporal dependencies and contribute to increased complexity Freire (2021). Analyzing the intricacies of NN architectures, including their depth, the number of parameters, and computational demands, is crucial for understanding the trade-offs between performance and complexity. Deep learning models often exhibit enhanced capabilities in capturing intricate patterns and relationships in optical communication data. Still, their complexity may pose challenges regarding training time, computational resources, and practical implementation Asghar (2019). A thorough examination of these complexities is essential for identifying optimal models, such as choosing between a CNN for image-based tasks or an LSTM for sequential data, that balance high performance and manageable complexity, facilitating their efficient integration into short-reach optical communication systems Freire (2021). Four prominent types, namely feed-forward neural networks (FFNN), radial basis function neural networks (RBF-NN), auto-regressive recurrent neural networks (AR-RNN), and layer-recurrent neural networks (L-RNN), offer distinct trade-offs in complexity and performance. Numerical evaluations reveal that, for a fixed bit error ratio (BER) threshold, AR-RNN-based equalizers exhibit the lowest computational complexity. Among nonlinear NN-based equalizers with equivalent numbers of inputs and hidden neurons, FFNN-based equalizers generally boast the lowest computational complexity. In contrast, AR-RNN-based equalizers demonstrate superior BER performance Xu (2019).
@@
Distillation model: Knowledge distillation, a model compression technique, transfers knowledge from intricate, large-scale models or ensembles to more compact, feasible models suitable for real-world constraints. Pioneered by Bucilua and collaborators in 2006 Buciluǎ (2006), knowledge distillation primarily operates on neural network architectures characterized by multifaceted structures comprising multiple layers and parameters. The surge of deep learning over the past decade, notably successful in domains like speech and image recognition and natural language processing Gou (2020), has propelled knowledge distillation as a pivotal methodology for practical applications. The demand for deploying extensive deep neural network models becomes particularly challenging for edge devices limited by memory and computational resources. In response, an innovative model compression method Gou (2020) emerged, facilitating the seamless transfer of knowledge from expansive models to train more resource-efficient counterparts without substantial performance degradation. This transformable process, where a smaller model acquires knowledge from a larger model, was systematized into the formal "Knowledge Distillation" framework by Hinton and colleagues Hinton (2015). This framework has become integral for addressing the computational constraints of edge devices while preserving the essential knowledge distilled from sophisticated, larger-scale models.
@@
Optimizing deep learning models through knowledge distillation holds substantial promise for future advancements in short-reach optical communication systems. Notably, recurrent neural networks (RNNs) have demonstrated unparalleled efficacy in equalizing nonlinear impairments Freire (2021) StavrosD (2021). However, the inherent feedback loop in RNN structures poses a formidable challenge to parallelization, hindering their implementation in low-complexity hardware tailored for high-speed processing in optical networks AndreXianMingC (2017). To overcome this limitation, a prospective avenue involves harnessing knowledge distillation as a strategic solution for enabling parallelization in RNNs Freire (2021) SasipimS (2023). This application of knowledge distillation is poised to revolutionize the implementation of RNNs, ensuring compatibility with low-complexity hardware and meeting the stringent processing requirements of high-speed optical networks RobertR (2021).
Looking beyond RNNs, the future trajectory of knowledge distillation extends its influence across various machine-learning models crucial to optical communications. Examples include CNNs, LSTMs, FFNN, RBF-NN, AR-RNN, and L-RNN SasipimS (2023). These diverse models present distinct challenges in complexity, adaptability, and real-time processing. For instance, consider the scenario of implementing knowledge distillation in CNNs designed for optical signal processing tasks. The distillation process can facilitate knowledge transfer from a more complex teacher model to a simplified student model, optimizing the trade-off between accuracy and computational efficiency. Similarly, in LSTMs utilized for sequence processing in optical communication systems, knowledge distillation can streamline the model complexity without compromising its ability to capture intricate temporal dependencies SasipimS (2023).
@@
Another field that holds promise in short-reach optical communication systems is knowledge distillation techniques that address challenges with limited time-series data. As "big data" impacts various fields, the scarcity of target events or high acquisition costs hinders machine learning tasks in specific scenarios. The proposed method, rooted in generalized distillation, utilizes "privileged information" from middle-time data during training. This novel approach enhances long-term predictions for small time-series datasets. Applied to optical communications, it offers a solution to data constraints, demonstrating efficacy on synthetic and real-world data. This distillation model presents a pathway for more robust predictive modeling in optical communication systems ShogoH (2019).
@@
Vector quantization: Vector quantization (VQ) is a model compression technique to mitigate the complexity of large-scale machine learning models. The essence of VQ lies in its ability to represent intricate data with a compact set of prototype vectors, thus significantly reducing the computational load during inference. This technique finds applications where limited resources require a balance between model efficiency and performance. The working principle of VQ involves partitioning the input space into regions, each associated with a representative prototype vector. During the encoding phase, input vectors are assigned to the nearest prototype, effectively quantizing the data. In contrast, the given prototype vectors are employed to reconstruct the original data during the decoding or reconstruction phase.
@@
The efficiency of VQ hinges on the strategic selection and adaptation of prototype vectors, optimizing their representation to encapsulate the essential information within the dataset RobertG (1984). By clustering and quantizing input vectors into a representative codebook, VQ enables encoding information in a more compact form. This is particularly beneficial in scenarios with limited data availability or high computational demands. For instance, consider the application of VQ in knowledge distillation for recurrent neural networks (RNNs), which face challenges in parallelization due to their feedback loop structure. Implementing VQ in the distillation process can help mitigate the parallelization issue by representing the essential information with a reduced set of prototype vectors XinW (2019) KashifR (2022). This compression not only aids in overcoming hardware complexity but also contributes to faster processing in high-speed optical networks.
@@
VQ typically involves an iterative process that refines the prototype vectors to enhance their representative capacity. This is commonly achieved through algorithms like the k-means clustering algorithm, where prototypes are adjusted to minimize the overall distortion or reconstruction error between the original data and the quantized representation. The iterative refinement of prototype vectors allows vector quantization to adapt to the inherent patterns and structures within the data, optimizing its compression capabilities while preserving critical information for training tasks RobertG (1984).
@@
Furthermore, VQ can be crucial in optimizing other deep learning models, such as CNNs or LSTMs, by efficiently capturing essential features with a minimal set of representative vectors SherjilO (2021). The future exploration of VQ in conjunction with these models offers a promising avenue for improving the performance and scalability of machine learning applications in short-reach optical communication systems.

### v2

@@ 5.3 Transformer-based Network
Attention: The Scaled Dot-Product Attention mechanism is the key component aiming to aggregate information across different parts of the input sequence. Each input vector is transformed into three distinct vectors: Queries ($Q$), Keys ($K$), and Values ($V$). The process involves calculating the dot products of the queries with all keys, scaling them by the square root of the dimension, and applying a softmax function to obtain weights on the values VaswaniA (2017).
@@
Transformers use multiple attention heads to look at the input sequence from different perspectives. This allows the model to simultaneously learn and consider various views of the input data. Equation 14 and 15 represent the functionality of the multi-head attention step. $head_{i}$ represents the single attention head. The final result of the multi-head attention is concatenating all the attention heads.
@@
While the standard (’vanilla’) transformer model has shown great performance on time series data, the computational complexity of its attention mechanism makes it struggle to handle long sequences effectively. To overcome this limitation, researchers have developed various attention mechanism variants. An example is the Locality-Sensitive Hashing (LSH) Attention mechanism, which was introduced as part of the Reformer architecture KitaevN (2020). The LSH Attention mechanism utilizes specialized hash functions to transform queries and keys, thereby facilitating the categorization of similar items into shared hash buckets. Through sorting tokens based on their hash codes, items with similarities are grouped together, enabling the aggregation of relevant information. To enable parallel processing, the sorted sequence is divided into chunks. Subsequently, attention mechanisms are selectively applied to these chunks and their neighboring segments, allowing for focused examination of localized portions. The LSH Attention mechanism uses hash coding to greatly improve the computational efficiency of the transformer compared to the original version. This helps address the challenge of processing long sequences by reducing complexity without sacrificing performance.
@@
Decoder: The decoder consists of several stacked sub-decoders. The ground truth follows a process similar to that of the encoder, being transformed into Query $Q$, Key $K$, and Value $V$ representations. The attention weights are calculated by comparing the Queries $Q$ from the decoder with every value $V$ in the encoder. This process is repeated in parallel across N sub-decoders, resulting in a final attention matrix. The attention matrix then undergoes a softmax operation, yielding probabilities for each value. In addition to the two sub-layers in each encoder layer, the decoder introduces a third sub-layer, performing multi-head attention over the encoder stack’s output.
@@
During the decoding process, the model is auto-regressive, using the previously generated outputs as additional input to generate the next output. Residual connections and layer normalization are used around each sub-layer. The self-attention sub-layer is modified to prevent positions from attending to future positions. This, along with the offset output embedding, ensures that the predictions for a position only depend on the known outputs at earlier positions in the sequence.
@@
Attention Variants: The purpose of the attention layer is to identify connections and dependencies among the various input embeddings. This allows the model to evaluate the importance of each element in relation to the others. The attention mechanism explicitly computes the relationships between different elements in the sequence, providing insights into how information flows through the model. However, except for the computational complexity issue, the mechanism in vanilla transformer VaswaniA (2017) needs to be improved in terms of processing inter-dependencies and periodicity of signal data. The Autoformer model WuH (2021) introduces a new type of encoder that replaces the original encoder. This new encoder applies series decomposition and autocorrelation to detect dependencies between different parts of the input sequence, and then combines the representations of the sub-series. The series decomposition component divides the original signal into two distinct parts: the seasonal component, which captures short-term patterns, and the trend component, which captures long-term behavior. This partitioning allows for identifying and representing both the short-term and long-term characteristics present in the time series data. Additionally, the auto-correlation mechanism utilizes the fast Fourier transform (FFT) to compute correlations between the time series and its delayed version, providing insights into how the series relates to its past values at different time lags. The combination of series decomposition and autocorrelation effectively captures and represents the underlying trend and seasonality in the time series data.
@@
Recent Progress: In this section, we provide a comprehensive overview of the vanilla transformer and its architecture, particularly within the domain of time series and traffic prediction. Over the years, substantial improvements have been made to enhance the transformer for accurate time series prediction. Notably, advancements have been achieved in reducing computational complexity while improving the effectiveness of the attention mechanism logsparse (2019); informer (2021). However, recent research has introduced compact models based on multi-scale transformation scaleformer (2023), which surprisingly outperforms benchmark-designed models. This new development has sparked an important debate on the fundamental structure of sequence models. In the following sections, we summarize and explore this particular model in depth, providing insights into its implications. For the latest work, please refer to Table 3 (table 3).
@@ 5.4 Fourier Convolution Neural Network
In the previous section, we introduced models based on convolutional kernels and subsampling as fundamental modules. Their core principle involves decomposing signals into different scales in the time domain and subsequently applying nonlinear transformations to learn salient features. However, for the majority of real-world signals, transforming them into the Fourier domain is often more efficient. This efficiency is attributed to the following factors: 1) The majority of real-world signals are bandpass or lowpass, and in the Fourier domain, their dynamic range decreases from $n$ to $exp(-n)$; 2) The Fourier transformation is a bijective (one-to-one) transformation, which ensures energy conservation and controllable error in both the forward and inverse transformations; 3) The computational complexity of existing (fast) Fourier transform algorithms, after improvements, is $O(nlog(n))$, making it convenient for hardware implementation.
In this section, we introduce TimesNet WuH (2022) which utilizes a frequency-attention mechanism and FreTS YiK (2023) which explicitly performs non-linear transformations in the frequency domain. For extensive models please refer to Table 3. The fundamental concept of TimesNet involves transforming the initial signal into $k$ distinct 2D tensors instead of directly processing the original sequence. This approach empowers the model to effectively capture both intra-periodic and inter-periodic variations within these fixed windows. A variant of this model has been recently reported in ShaoFConv (2024)
@@
Component Detection: The identification of the k most crucial frequencies is based on the amplitude of the Fourier coefficients. Then, using only the selected components within the k frequency range, the signal is sampled using a continuous sampling method, and these sub-series are arranged into a two-dimensional tensor.
Alignment: The aligned sub-series are then fed into a convolution-based module, specifically an Inception network, to mitigate distortion caused by channel effects. The Fourier coefficients pass through a softmax function to generate attention weights, which are then multiplied by the output of each convolution module to produce the final output.
Fourier Attention: The Fourier transformation is a global operation, meaning that any changes in the signal’s amplitude will cause periodic oscillations throughout the entire signal. Significant variations in the amplitude of the primary components lead to substantial fluctuations. TimesNet leverages this characteristic by using the Fourier spectrum of the nonlinearly transformed signal to determine the attention value for each component.
@@
FreTS: Time-domain-based processing is limited by information bottlenecks, as the local characteristics vary. FreTS explicitly uses frequency-domain features in its model architecture to directly mitigate distortion without manipulating the time-domain. FreTS is essentially an MLP-based network that is able to effectively learn patterns of time series data in the frequency domain. As presented in YiK (2023), FreTS consists of two learners: a Frequency Channel Learner and a Frequency Temporal Learner. In the equalization problem, there is no actual channel dimension, but rather a stack of independent experiments. Therefore, FreTS only introduces a frequency domain MLP.
Frequency MLP: The frequency temporal learner aims to capture temporal patterns in the frequency domain. Specifically, for a complex number input $\mathcal{H}\in\mathbb{C}^{m\times d}$, the MLP aims to optimize the weight matrix $\mathcal{W}\in\mathbb{C}^{d\times d}$ and bias $\mathcal{B}\in\mathbb{C}^{d}$ so that the final output $\mathcal{Y}\in\mathbb{C}^{m\times d}$ could approximately reconstruct the ground truth.
@@
The MLP in the frequency domain is equivalent to global convolutions in the time domain as detailed in YiK (2023). An increasing number of studies have demonstrated the feasibility of DL models operating in the frequency domain. Simultaneously, the corresponding computational complexity of frequency-domain processing has decreased from $O(n)$ to $O(nlogn)$ due to the reduction in the signal’s dynamic range. However, the advantages and disadvantages of networks in both the frequency and time domains remain inadequately explored. Due to space limitations, we offer a detailed categorization, along with corresponding references and keywords, in Table 3 for researchers with specific interests.
@@ 6 Model Compression
In recent years, the proliferation of large-scale ML models has significantly advanced state-of-the-art technology across various domains, ranging from natural language processing to computer vision. The surge in model complexity, often characterized by sophisticated architectures and many parameters, has driven the need for efficient hardware implementations to harness their full potential. The advent of single graphics processing units (GPUs) as a critical computational resource has been pivotal, offering a parallelized architecture suitable for accelerating the training and inference processes LewJ (2019). The significance of deploying large ML models on a single GPU lies in optimizing computational efficiency and reducing latency. Single GPU implementations facilitate parallel processing, enabling the simultaneous execution of multiple tasks and handling extensive model parameters. This enhances the speed of model training and facilitates real-time inference, a critical requirement in applications such as autonomous systems and edge computing. However, the hardware implementation of large ML models on a single GPU is challenging. The complexity of these models often exceeds the computational capacity and memory constraints of a single GPU, necessitating innovative solutions for efficient utilization MarculescuD (2018). Techniques such as model pruning, quantization, vector quantization, and knowledge distillation have emerged as strategies to mitigate these challenges, ensuring that even formidable models can be accommodated within the limitations of a single GPU without compromising performance. The authors in MarculescuD (2018) examine how to use a single GPU effectively for implementing large ML models. They discuss methods that balance complexity and computational efficiency to maximize hardware utilization MarculescuD (2018).
@@
In addition, conducting a comprehensive performance-versus-complexity analysis is necessary to evaluate the suitability of various ANNs in short-reach optical communication systems. DL models, including CNNs, RNNs, and LSTMs, find applications in critical tasks such as equalization, fault detection, subcarrier allocation, nonlinearity compensation, and bandwidth request and allocation. The complexity of these models is a significant factor affecting their feasibility. For instance, CNNs may introduce convolutional and pooling layers, increasing model complexity. Similarly, RNNs and LSTMs, designed for sequential data, introduce recurrent connections that enhance their ability to capture temporal dependencies and contribute to increased complexity Freire (2021). Analyzing the neural network architectures in detail, including their depth, the number of parameters, and computational demands, is crucial for understanding the trade-offs between performance and complexity. DL models often exhibit enhanced capabilities in capturing complex patterns and relationships in optical communication data. Still, their complexity may pose challenges regarding training time, computational resources, and practical implementation Asghar (2019). A thorough examination of these complexities is essential for identifying optimal models, such as choosing between a CNN for image-based tasks or an LSTM for sequential data, that balance high performance and manageable complexity, facilitating their efficient integration into short-reach optical communication systems Freire (2021). Four prominent types, namely the feed-forward neural networks (FFNN), the radial basis function neural networks (RBF-NN), the auto-regressive RNN (AR-RNN), and layer-RNN (L-RNN), offer distinct trade-offs in complexity and performance. Among nonlinear neural network-based equalizers with equivalent numbers of inputs and hidden neurons, FFNN-based equalizers have the lowest computational complexity, however AR-RNN demonstrate superior transmission performance in 50 Gb/s PAM4 systems Xu (2019).
@@
Distillation model: Knowledge distillation, a model compression technique, transfers knowledge from complex, large-scale models or groups to more compact, feasible models suitable for real-world applications. Pioneered by Bucilua et al. in 2006 Buciluǎ (2006), knowledge distillation primarily operates on neural network architectures characterized by multifaceted structures comprising multiple layers and parameters. Knowledge distillation has been recently considered an important technique for practical DL applications such as speech recognition, image recognition, and natural language processing Gou (2020). Deploying large deep neural network models can be especially challenging for edge devices, which are limited in memory and computational power. To address this challenge, an innovative model compression method was developed in Gou (2020), allowing transferring knowledge from larger, more complex models to train smaller, more efficient models without significant performance loss. This process, where a smaller model learns from a larger one, was formalized into the "Knowledge Distillation" framework by Hinton et al. Hinton (2015). This framework has become crucial for deploying the essential knowledge from sophisticated, large-scale models on computationally constrained edge devices.
@@
Optimizing DL models through knowledge distillation shows great potential for advancing short-range optical communication systems. RNNs have been particularly effective at addressing nonlinear distortions Freire (2021); StavrosD (2021). However, the feedback loop inherent in RNN structures makes it difficult to parallelize them, preventing their use in low-complexity hardware designed for high-speed processing in optical networks AndreXianMingC (2017). Using knowledge distillation is a promising approach to enable parallelization of RNNs Freire (2021); SasipimS (2023). This application of knowledge distillation is set to revolutionize the implementation of RNNs, ensuring compatibility with low-complexity hardware and meeting the stringent processing requirements of high-speed optical networks RobertR (2021).
Beyond just RNNs, knowledge distillation can be applied to many different ML models important for optical communications, such as CNNs, LSTMs, FFNNs, RBF-NNs, AR-RNNs, and L-RNNs SasipimS (2023). These models each have their own challenges regarding complexity, adaptability, and real-time implementation. For example, using knowledge distillation in LSTMs for optical communication systems, can reduce model complexity without losing the ability to handle time-dependent patterns SasipimS (2023).
@@
Another promising application of knowledge distillation is when facing challenges with limited time-series data. As "big data" impacts various fields, the scarcity of target events or high data acquisition costs can hinder ML in certain scenarios. A proposed method uses "privileged information" from partial time-series data during training to enhance long-term predictions for small datasets. Applied to optical communications, this distillation approach offers a solution to data constraints, demonstrating effectiveness on both synthetic and real-world data ShogoH (2019).
@@
Vector quantization: Vector quantization (VQ) is a model compression technique targeting large-scale ML models. VQ represents complex data with a small set of prototype vectors, significantly cutting the computational load during inference. This makes VQ useful for applications that require balanced model efficiency and performance, such as when resources are limited. The VQ process involves partitioning the input space into regions, each with a representative prototype vector. During encoding, input vectors are assigned to the nearest prototype, quantizing the data. In the decoding or reconstruction phase, these prototype vectors are used to rebuild the original data.
@@
The effectiveness of VQ relies on carefully selecting and updating the prototype vectors. The goal is to optimize the prototypes so they can effectively capture the essential information in the dataset RobertG (1984). By clustering and quantizing input vectors into a representative codebook, VQ enables encoding information in a more compact form. This is particularly beneficial in scenarios with limited data availability or high computational demands HamidRezaP (2020). For example, VQ can be useful when applying knowledge distillation to RNNs. RNNs face challenges with parallelization due to their feedback loop structure. Using VQ in the distillation process for RNNs can help address the parallelization issue. VQ can represent the essential information from the RNN using a smaller set of prototype vectors. Xu (2019); XinW (2019); KashifR (2022). This compression not only aids in overcoming hardware complexity but also contributes to faster processing in high-speed optical networks.
@@
VQ uses an iterative process to improve the prototype vectors and enhance their ability to represent the data. Commonly, algorithms like k-means clustering are used for this. The prototypes are adjusted to minimize the difference between the original data and the quantized representation. This iterative refinement allows VQ to adapt to the patterns and structures in the data. This optimizes the compression capabilities of VQ while still preserving the critical information needed for training tasks RobertG (1984).
@@
Finally, VQ can be beneficial in optimizing other DL models, such as CNNs or LSTMs, by efficiently capturing essential features with a minimal set of representative vectors SherjilO (2021). Exploring the use of VQ together with these models provides a promising way to improve the performance and scalability of ML applications in short-reach optical communication systems.